# Representation Learning for Distributional Perturbation Extrapolation

### v1

@@ B.4 Perturbation Modelling
$\widehat{\mathbf{z}}^{\mathrm{base}}=\mathbf{g}(\mathbf{x})\,.$ (B.14)
@@
The estimated perturbed latent state $\widehat{\mathbf{z}}^{\mathrm{pert}}$ is then computed according to (B.12) using (B.14) and learnt estimates $\widehat{\mathbf{W}}^{\mathrm{pert}}$, $\{\widehat{h}_{k}\}$, and $\{\widehat{\mathbf{W}}_{l}^{\text{cov}}\}$. Finally, a (deterministic) decoder $\mathbf{f}$ uses $\widehat{\mathbf{z}}^{\mathrm{pert}}$ to compute estimates of the mean and variance in (B.13), i.e., $(\widehat{\bm{\mu}},\widehat{\sigma}^{2})=\mathbf{f}(\widehat{\mathbf{z}}^{\mathrm{pert}})$. All learnable components of the model are trained by minimizing the (Gaussian) negative log-likelihood of the observed data $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{l}_{i},\mathbf{C}_{i})\}_{i\in[N]}$. To encourage the postulated independence of $\widehat{\mathbf{z}}^{\mathrm{base}}$ and $(\mathbf{l},\mathbf{C})$, an additional adversarial loss is used, which minimizes the predictability of the latter from the former.
Appendix C Relation to Causal Models
In the field of causal inference, experimental data resulting from perturbations is typically modelled via interventions in an underlying causal model. In the structural causal model (SCM) framework of Pearl (2009), interventions modify a set of assignments, which determine each variable as a function of its direct causes and unexplained noise.
Definition C.1 (Acyclic SCM).
An acyclic structural causal model (SCM) $\mathcal{M}=\left({\mathcal{S}},\mathbb{P}_{\mathbf{U}}\right)$ with endogenous variables $\mathbf{V}=\{V_{1},...,V_{n}\}$ and exogenous variables $\mathbf{U}=\{U_{1},...,U_{m}\}$ consists of:
  1. (i)
    a set of structural equations
    ${\mathcal{S}}=\left\{V_{i}:=f_{i}\left(\mathbf{V}_{\mathrm{pa}(i)},\mathbf{U}_{i}\right)\right\}_{i=1}^{n},$ (C.1)
    where $f_{i}$ are deterministic functions; $\mathbf{U}_{i}\subseteq\mathbf{U}$; and $\mathbf{V}_{\mathrm{pa}(i)}\subseteq\mathbf{V}\setminus\{V_{i}\}$ is the set of causal parents of $V_{i}$, such that the induced causal graph with vertices $[n]$ and edges $j\to i$ iff. $j\in\mathrm{pa}(i)$ is acyclic;
  2. (ii)
    a joint distribution $\mathbb{P}_{\mathbf{U}}$ over the exogenous variables.
The induced distribution $\mathbb{P}_{\mathbf{V}}$ of $\mathcal{M}$ is given by the push-forward of $\mathbb{P}_{\mathbf{U}}$ via ${\mathcal{S}}$. An intervention replaces ${\mathcal{S}}$ by a new set of assignments ${\mathcal{S}}^{\prime}$ such the graph induced by ${\mathcal{S}}^{\prime}$ is acyclic. The interventional distribution is the induced distribution $\mathcal{M}^{\prime}=({\mathcal{S}}^{\prime},\mathbb{P}_{\mathbf{U}})$. [^7] Interventions can also introduce new sources of randomness $\mathbf{U}^{\prime}$, which, for sake of simplicity, we do not consider here.
Our assumed generative process for the effect of perturbations (§ 3) is, in general, different from interventions in an SCM. However, as we show next, certain types of SCMs and interventions are recovered as a special case of our model.
Shift Interventions in Linear SCMs as a Special Case.
Consider a linear SCM with ${d_{Z}}$ endogenous variables $\mathbf{Z}$ and ${d_{Z}}$ exogenous variables $\mathbf{U}$ of the form
$\mathbf{Z}:=\mathbf{A}\mathbf{Z}+\mathbf{U},$ (C.2)
where $\mathbf{A}$ is a (lower-triangular) weighted adjacency matrix. The observational (i.e., non-intervened) distribution of $\mathbf{Z}$ induced by (C.2) is most easily understood via the reduced form expression
$\mathbf{Z}=(\mathbf{I}-\mathbf{A})^{-1}\mathbf{U}$ (C.3)
and thus given by $\mathbb{P}_{\mathbf{Z}}=(\mathbf{I}-\mathbf{A})^{-1}_{\#}\mathbb{P}_{\mathbf{U}}$. Now consider the class of shift interventions parametrised by constant shift vectors $\mathbf{c}_{e}\in\mathbb{R}^{{d_{Z}}}$, which modify the original SCM in (C.2) to
$\mathbf{Z}_{e}:=\mathbf{A}\mathbf{Z}_{e}+\mathbf{U}+\mathbf{c}_{e}.$ (C.4)
Analogous to (C.3), the reduced form of (C.4) is given by
$\mathbf{Z}_{e}$ $=(\mathbf{I}-\mathbf{A})^{-1}\mathbf{U}$ $+(\mathbf{I}-\mathbf{A})^{-1}\mathbf{c}_{e}$ (C.5)
$=\qquad\mathbf{Z}$ $+(\mathbf{I}-\mathbf{A})^{-1}\mathbf{c}_{e},$
where the second equality follows from (C.3).
Thus, if we take the shift vectors as perturbation labels (i.e., $K={d_{Z}}$ and $\mathbf{l}_{e}=\mathbf{c}_{e})$ and use a linear perturbation model of the form
$\mathbf{Z}^{\mathrm{pert}}_{e}=\bm{\phi}(\mathbf{Z}^{\mathrm{base}}_{e},\mathbf{l}_{e})=\mathbf{Z}^{\mathrm{base}}_{e}+\mathbf{W}\mathbf{l}_{e}$ (C.6)
with $\mathbf{W}=(\mathbf{I}-\mathbf{A})^{-1}$, then our perturbation model captures shift interventions in a linear SCM with adjacency matrix $\mathbf{A}$.
Remark C.2.
The above argument does not require causal sufficiency: it still holds if $\mathbb{P}_{\mathbf{U}}$ is not factorized (e.g., due to hidden confounding).
Remark C.3.
When trained on data generated according to (C.4), the adjacency matrix associated with the learnt perturbation matrix $\widehat{\mathbf{W}}$ is given by $\widehat{\mathbf{A}}=\mathbf{I}-\widehat{\mathbf{W}}^{-1}$.
Remark C.4.
The correspondence between mean shift perturbations and shift interventions appears to only hold for linear SCMs. Consider, for example, a nonlinear additive noise model,
$\mathbf{Z}:=\mathbf{f}(\mathbf{Z})+\mathbf{U},$ (C.7)
with reduced form given by
$\mathbf{Z}=\mathbf{g}(\mathbf{U}),$ (C.8)
where $\mathbf{g}$ is the inverse of the mapping $\mathbf{z}\mapsto\mathbf{z}-\mathbf{f}(\mathbf{z})$. For shift interventions in (C.7) to match our perturbation model, we then must have
$\mathbf{Z}_{e}=\mathbf{g}(\mathbf{U}+\mathbf{c}_{e})=\mathbf{g}\left(\mathbf{g}^{-1}(\mathbf{Z})+\mathbf{c}_{e}\right)=\bm{\phi}\left(\mathbf{Z},\mathbf{l}_{e}\right)$ (C.9)
for suitable $\bm{\phi}$ and $\mathbf{l}_{e}$. Thus, if $\mathbf{f}$ is nonlinear, so is $\mathbf{g}$ and therefore $\bm{\phi}$. In other words, shift interventions in a nonlinear SCM do not, in general, amount to mean shift perturbations (i.e., linear $\bm{\phi}$).
Appendix D Additional Experimental Details
Complete List of Test Cases.
The six ID and OOD test cases used for evaluation in the experiments from § 7 are listed in Tab. 2. The resulting ground truth distributions for all synthetic train and test distributions are visualized in Fig. 4.
Table 2: Test Perturbation Labels $\mathbf{l}_{\mathrm{test}}^{\mathrm{ID}}$ and $\mathbf{l}_{\mathrm{test}}^{\mathrm{OOD}}$ for the Experiment on 2D Data From § 7.
ID OOD
[1, 1, 0]11 0[1,\ 1,\ 0][ 1 , 1 , 0 ] [1, 0, 1]1 01[1,\ 0,\ 1][ 1 , 0 , 1 ]
[0.1, 0.25, 0]0.10.25 0[0.1,\ 0.25,\ 0][ 0.1 , 0.25 , 0 ] [0, 1, 1]011[0,\ 1,\ 1][ 0 , 1 , 1 ]
[0.5, 0.75, 0]0.50.75 0[0.5,\ 0.75,\ 0][ 0.5 , 0.75 , 0 ] [0, 2, 0]02 0[0,\ 2,\ 0][ 0 , 2 , 0 ]
[0.5, 0.5, 0]0.50.5 0[0.5,\ 0.5,\ 0][ 0.5 , 0.5 , 0 ] [2, 0, 0]2 0 0[2,\ 0,\ 0][ 2 , 0 , 0 ]
[0, 0.5, 0]00.5 0[0,\ 0.5,\ 0][ 0 , 0.5 , 0 ] [0.5, 1.5, 0]0.51.5 0[0.5,\ 1.5,\ 0][ 0.5 , 1.5 , 0 ]
[0.5, 0, 0]0.5 0 0[0.5,\ 0,\ 0][ 0.5 , 0 , 0 ] [1.5, 0.5, 0]1.50.5 0[1.5,\ 0.5,\ 0][ 1.5 , 0.5 , 0 ]
Figure 4: Complete Ground Truth Training and Test Data. Shown are all train and test distributions, with crosses marking the respective means. The training cases are depicted in grey, the ID test cases in blue and the OOD test cases in orange.
Hyperparameter Selection for CPA.
To improve the fit of CPA on the synthetic dataset, we perform a grid search over key hyperparameters. Specifically, we consider learning rates of the autoencoder and adversarial classifier from {1e-2, 5e-3, 1e-3, 5e-4, 1e-4, 5e-5}, and a latent bottleneck dimension from $\hat{d_{Z}}$ in {2, 4, 8}. We select the best joint parameter configuration in terms of a combined performance measure (ED + MMD + mean difference) on a held out portion of 25% of the original training data, which yields a learning rate of 1e-3 for the autoencoder, a learning rate of 5e-3 for the adversarial classifier, and a bottleneck dimension of 4. All other hyperparameters are set to their default values.
Appendix E Additional Results
In this experiment, each observation $\mathbf{x}_{e,i}$ is generated from two signal dimensions $\mathbf{z}_{e,i}^{\text{pert}}\in\mathbb{R}^{2}$ and eight noise dimensions $\varepsilon_{e,i}^{\text{pert}}\in\mathbb{R}^{8}$ as input for the nonlinear mixing function, as described at the end of § 7. We evaluate model robustness under increasing levels of noise by varying the standard deviation $\sigma_{\varepsilon}$ of the noise variable $\varepsilon\sim\mathbb{Q}_{\epsilon}$ in the data-generating process. The list of values of $\sigma_{\epsilon}$ considered is [0, 0.01, 0.1, 0.25, 0.5, 1, 2, 10]00.010.10.250.51210[0,\ 0.01,\ 0.1,\ 0.25,\ 0.5,\ 1,\ 2,\ 10][ 0 , 0.01 , 0.1 , 0.25 , 0.5 , 1 , 2 , 10 ]. Both PDAE and CPA models are trained for 2000 epochs. One random seed is used and shared across data generation, model initialization, training, and evaluation. We use the same hyperparameters as described in § 7, with the exception of PDAE, for which we increase the reconstruction loss weight $\lambda=0.25$ and set the decoder noise dimension to 1 with a standard deviation of 0.5; and CPA, for which we use a 2-dimensional embedding. The results are summarised in Fig. 5. As can be seen, PDAE performs competitively across a range of $\sigma_{\epsilon}$ (especially low to moderate ones), despite only having access to one noise dimension in the decoder. As $\sigma_{\epsilon}$ grows large, most signal is lost, and the performance of all methods tends to that of the pooling baseline, as expected.

### v2

@@ C.4 Perturbation modeling
$\widehat{\mathbf{z}}^{\mathrm{base}}=\mathbf{g}(\mathbf{x})\,.$ (C.14)
@@
The estimated perturbed latent state $\widehat{\mathbf{z}}^{\mathrm{pert}}$ is then computed according to (C.12) using (C.14) and learnt estimates $\widehat{\mathbf{W}}^{\mathrm{pert}}$, $\{\widehat{h}_{k}\}$, and $\{\widehat{\mathbf{W}}_{l}^{\text{cov}}\}$. Finally, a (deterministic) decoder $\mathbf{f}$ uses $\widehat{\mathbf{z}}^{\mathrm{pert}}$ to compute estimates of the mean and variance in (C.13), i.e., $(\widehat{\bm{\mu}},\widehat{\sigma}^{2})=\mathbf{f}(\widehat{\mathbf{z}}^{\mathrm{pert}})$. All learnable components of the model are trained by minimizing the (Gaussian) negative log-likelihood of the observed data $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{a}_{i},\mathbf{C}_{i})\}_{i\in[N]}$. To encourage the postulated independence of $\widehat{\mathbf{z}}^{\mathrm{base}}$ and $(\mathbf{a},\mathbf{C})$, an additional adversarial loss is used, which minimizes the predictability of the latter from the former.
Appendix D Additional experimental details and results
Figure 6: Qualitative comparison of PDAE and CPA on synthetic 2D data. Rows correspond to latent space (top) and observation space (bottom). Columns show the ground truth data (left), PDAE predictions (center), and CPA predictions (right). Training domains are shown in grey, an in-distribution (ID) test case with $\mathbf{a}^{\textsc{id}}_{\mathrm{test}}\!=\!(1,1,0)^{\top}$ (which overlaps with one of the training domains) in blue, and an out-of-distribution (OOD) test case with $\mathbf{a}_{\mathrm{test}}^{\textsc{ood}}\!=\!(1,0,1)^{\top}$ in orange. All plots show kernel density estimates of the distributions; crosses (x) indicate the corresponding means. As can be seen, PDAE recovers an affine transformation of the true latents (top, center) leading to accurate distributional predictions for the training and ID test domain (bottom, center). However, the OOD test domain is mapped to a part of the latent space not seen during training (top, center). As a result, the corresponding decoder output does not accurately match the true OOD distribution (bottom left vs center). CPA accurately predicts the means of the training distribution (bottom right, black crosses) but does not recover the true latents up to an affine transformation (top right). Moreover, the predicted distributions for CPA do not match the ground truth particularly well, particularly for the test conditions (bottom left vs right). However, recall that—unlike PDAE—CPA is not trained for distributional reconstruction, see §˜C.4 for details.
Table 2: Tuned Hyperparameters of CPA with their search space and optimal value. We sample 100 random configurations from the joint search space and train a CPA model for 2000 epochs for each. The optimal configuration minimizes the MMD on ID validation data.
Hyperparameter Search Space Selected Optimal Value
step_size_lr $\{15,\,25,\,45\}$ $15$
autoencoder_lr $10^{\mathcal{U}[-5,\,-2]}$ $3.63\text{\times}{10}^{-4}$
autoencoder_wd $10^{\mathcal{U}[-8,\,-4]}$ $5.97\text{\times}{10}^{-6}$
reg_adversary $10^{\mathcal{U}[-2,\,2]}$ $2.31\text{\times}{10}^{-2}$
penalty_adversary $10^{\mathcal{U}[-2,\,2]}$ $8.15\text{\times}{10}^{-2}$
adversary_lr $10^{\mathcal{U}[-5,\,-2]}$ $1.7\text{\times}{10}^{-4}$
adversary_wd $10^{\mathcal{U}[-8,\,-4]}$ $3.11\text{\times}{10}^{-6}$
adversary_steps $\{1,\,2,\,3,\,4,\,5\}$ 55
D.1 Hyperparameter selection for CPA
For CPA, hyperparameters are chosen by randomly searching over a predefined parameter space for 100 trials for every dataset, see Tab. 1 of Lotfollahi et al. (2023) for details. To improve the CPA fit on our simulation data and render the comparison with PDAE fairer, we mimic this procedure and run a random search for 100 trials over a subset of these hyperparameters, which we show in Tab.˜2. We tune all parameters that are neither linked to the architecture nor related to the non-linear dose-response functions. (In the setting of gene perturbation experiments that we emulate here, the dose responses are set to the identity, and thus no doser MLP hyperparameters are needed.) After randomly choosing 100 configurations of the parameter space and training a CPA model for 2000 epochs on each, we compute the MMD between predictions and ground truth of 7 ID evaluation environments and select the configuration that minimizes the MMD. The optimal configuration is shown in Tab.˜2.
@@ D.1 Hyperparameter selection for CPA
D.2 Robustness experiment with varying signal-to-noise ratio
In this experiment, each observation $\mathbf{x}_{e,i}$ is generated from two signal dimensions $\mathbf{z}_{e,i}^{\text{pert}}\in\mathbb{R}^{2}$ and eight noise dimensions $\varepsilon_{e,i}^{\text{pert}}\in\mathbb{R}^{8}$ as input for the nonlinear mixing function, as described at the end of §˜7. We evaluate model robustness under increasing levels of noise by varying the standard deviation $\sigma_{\varepsilon}$ of the noise variable $\varepsilon\sim\mathbb{Q}_{\epsilon}$ in the data-generating process. The list of values of $\sigma_{\epsilon}$ considered is [0, 0.1, 0.25, 0.5, 1, 2, 10][0,\ 0.1,\ 0.25,\ 0.5,\ 1,\ 2,\ 10]. We choose the same hyperparameters as described in §˜7, with the exception of increasing CPA’s bottleneck dimension to 10 and PDAE’s decoder model noise dimension to 8. Like this, both models are equipped with a decoder input dimension that fits the dimensionality of observations. On each noise level, we run 8 trials of random search for both methods. For CPA, we tune the set of parameters described in §˜7. For PDAE, we tune the reconstruction loss weight $\lambda$, the level of decoder noise and the learning rate. For both methods, the model with the lowest MMD on 7 ID validation environments is selected. Both PDAE and CPA models are then trained for 2000 epochs and evaluated on 7 ID test environments. One random seed is used and shared across data generation, model initialization, training, and evaluation. The results are summarised in Fig.˜7. As can be seen, PDAE outperforms the other methods across low and moderate levels of $\sigma_{\epsilon}$ and converges to the performance of Linear Regression at higher levels of noise. As $\sigma_{\epsilon}$ grows large, most signal is lost, and the performance of all methods tends to that of the pooling baseline, as expected.
1212OODID
Figure 8: (a) Sampling space for ID and OOD test cases. Representation of the ID region (in blue) as defined by the elementary mean shifts in the perturbation matrix and a possible OOD region (in orange) for mean shifts in the Euclidean plane. (b) Mean shifts of generated ID and OOD perturbation labels. Mean shifts from single perturbations (left) are generated from 3 different sampling schemes as detailed above. Mean shifts from double perturbations (right) are generated from 4 different sampling schemes. Overall, (3+4)10(3+4)*10 shifts are visualized for ID and OOD.
D.3 Details on validation and test cases
For every random seed, we randomly generate 14 ID and OOD perturbation labels. From the fixed training perturbations and the random ID and OOD perturbations, we then generate the two-dimensional latents and corresponding observations. The 14 ID perturbation labels are then split into 7 validation labels used for model selection (hyperparameter tuning) and 7 test labels to evaluate performance. An example of the resulting ground truth distributions for all synthetic train and test cases is visualized in Fig.˜9.
Random sampling of test cases. To make the choice of perturbation labels for validation and testing purposes less arbitrary, we set up a random sampling scheme. We propose two separate schemes for generating ID and OOD perturbation labels since their corresponding mean shifts lie within disjoint regions of the Euclidean plane, as shown in Fig.˜8(a). The ID region of mean shifts is constrained by our training perturbation labels $\mathbf{a}_{0},...,\mathbf{a}_{3}$. Together with the elementary mean shifts $\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3}$, i.e., the columns of the perturbation matrix $\mathbf{W}$, the columns of $\mathbf{W}\mathbf{A}$ constitute the vertices of the unit square as depicted in Fig.˜8(a). For ease of visualization, we choose the square in the first quadrant with origin at $[0,0]$ and edge length of 2 to constrain the region of OOD mean shifts. We now explain how the ID and OOD perturbation labels are sampled such that the corresponding mean shifts adhere to the ID and OOD region, respectively.
Single perturbation labels like our training labels only contain one non-zero entry. We therefore sample uniformly from
$\mathcal{R}^{\text{ID}}_{\text{single}}=\Big\{(a,0,0):a\in[0,1]\Big\}\cup\Big\{(0,b,0):b\in[0,1]\Big\}\cup\Big\{(0,0,c):c\in[0,1]\Big\}$ (D.1)
For the OOD single perturbation labels, the the non-zero entries lie in $[1,2]$
$\mathcal{R}^{\text{OOD}}_{\text{single}}=\Big\{(a,0,0):a\in[1,2]\Big\}\cup\Big\{(0,b,0):b\in[1,2]\Big\}\cup\Big\{(0,0,c):c\in[1,2]\Big\}$ (D.2)
Similarly, to generate double perturbations from the ID region, we draw uniformly from
$\mathcal{R}^{\text{ID}}_{\text{double}}=\Big\{[a,b,0]:a,b\in[0,1]\Big\}\cup\Big\{[a,0,c]:c\in[0,1],a\in[0,1-c]\Big\}\cup\Big\{[0,b,c]:c\in[0,1],b\in[0,1-c]\Big\}$ (D.3)
where the restriction to [1c][1-c] ensures that shifts remain within the ID region.
To create OOD double perturbation labels that imply admissible mean shifts, we define
$\mathcal{R}_{1}$ $=\Big\{[a,b,0]:a\in[0,2],b_{0}\in[0,1],b=\bm{1}_{\{a\geq 1\}}2b_{0}+\bm{1}_{\{a<1\}}(b_{0}+1)\Big\}$ (D.4)
$\mathcal{R}_{2}$ $=\Big\{[a,b,0]:b\in[0,2],a_{0}\in[0,1],a=\bm{1}_{\{b\geq 1\}}2a_{0}+\bm{1}_{\{b<1\}}(a_{0}+1)\Big\}$ (D.5)
$\mathcal{R}_{3}$ $=\Big\{[a,0,c]:c\in[0,2],a_{0}\in[0,1],a=a_{0}(2-\max(1,c))+\max(1-c,0)\Big\}$ (D.6)
$\mathcal{R}_{4}$ $=\Big\{[0,b,c]:c\in[0,2],b_{0}\in[0,1],b=b_{0}(2-\max(1,c))+\max(1-c,0)\Big\}$ (D.7)
and then draw uniformly from
$\mathcal{R}^{\text{OOD}}_{\text{double}}=\mathcal{R}_{1}\cup\mathcal{R}_{2}\cup\mathcal{R}_{3}\cup\mathcal{R}_{4}.$ (D.8)
We visualize the effects of 70 generated ID and 70 OOD perturbation labels as mean shifts of base state latents in Fig.˜8(b), i.e., we visualize $\bm{\mu}_{e}^{z}=\mathbf{W}\mathbf{a}_{e}$ for $e\in\{1,...,140\}$.
The number of perturbation labels is chosen as a multiple of 7 since we construct 3 single and 4 double perturbation settings for ID and OOD.
The implied mean shifts for the 4 training, 7 generated ID validation, 7 generated ID test and 14 generated OOD test environments is visualized in Fig.˜9.
Figure 9: Fixed train environments and randomly sampled validation and test environments with continous perturbation labels We visualize the Kernel Density Estimates of the ground truth distribution of the 4 fixed training environments as well as the randomly sampled 14 ID and 14 OOD environments. The ID environments are split into 7 for validation and 7 for testing. The means of the distributions are visualized with crosses.